OpenStudy (anonymous):

Choose the correct simplification of (a5b)(a2b).

OpenStudy (theviper):

Can u draw it ?

OpenStudy (hba):

Is it, \[\huge\ (a^5b)(a^2b)\]

OpenStudy (anonymous):

yes

OpenStudy (hba):

If that is your question,See this...\[\huge\ a^x \times a^y=a^{x+y}\]

OpenStudy (anonymous):

dont understand ?

OpenStudy (hba):

So, a^5 x a^2 x b^1 x b^1

OpenStudy (hba):

Just add the powers.

OpenStudy (anonymous):

ok :)

OpenStudy (theviper):

\[(a^5b)(a^2b)\]\[\large{=a^{5+2}b+b^{1+1}=??}\]

OpenStudy (hba):

Well,Viper wanted to say, \[\huge\ a^{5+2} b^{1+1}\]

OpenStudy (anonymous):

a^7 b^2

OpenStudy (hba):

Yeah right :)

OpenStudy (theviper):

right :)

OpenStudy (theviper):

Another question @andrewsurfs4995 ???

OpenStudy (anonymous):

yess:)

OpenStudy (theviper):

Then ask {not here}:D